# Code Prompting Elicits Conditional Reasoning Abilities in Text+Code LLMs

### 18 Jan 2024 v1

@@ Abstract
Reasoning is a fundamental component for achieving language understanding. Among the multiple types of reasoning, conditional reasoning, the ability to draw different conclusions depending on some condition, has been understudied in large language models (LLMs). Recent prompting methods, such as chain of thought, have significantly improved LLMs on reasoning tasks. Nevertheless, there is still little understanding of what triggers reasoning abilities in LLMs. We hypothesize that code prompts can trigger conditional reasoning in LLMs trained on text and code. We propose a chain of prompts that transforms a natural language problem into code and prompts the LLM with the generated code. Our experiments find that code prompts exhibit a performance boost between $2.6$ and $7.7$ points on GPT 3.5 across multiple datasets requiring conditional reasoning. We then conduct experiments to discover how code prompts elicit conditional reasoning abilities and through which features. We observe that prompts need to contain natural language text accompanied by high-quality code that closely represents the semantics of the instance text. Furthermore, we show that code prompts are more efficient, requiring fewer demonstrations, and that they trigger superior state tracking of variables or key entities. [^1] Code, prompt templates, prompts, and outputs are publicly available at https://github.com/UKPLab/arxiv2024-conditional-reasoning-llms.
@@ 1 Introduction
Reasoning is a fundamental component of both human and artificial intelligence (AI). Some reasoning types have received much attention in the form of new benchmarks and models from the AI community in the last few years, such as mathematical (Patel et al., 2021; Chen et al., 2021; Cobbe et al., 2021), logical (Liu et al., 2020, 2023a; Sinha et al., 2019), and commonsense reasoning (Madaan et al., 2022; Liu et al., 2022a, b; Wang et al., 2023). However, other important reasoning types, such as conditional reasoning, remain understudied. Conditional reasoning draws alternative conclusions depending on the fulfillment of certain conditions. It is also a fundamental form of logical reasoning useful in many scenarios, such as in chatbots (e.g., ChatGPT), to answer real-world questions like eligibility for a visa or a loan. Despite the recent introduction of some benchmarks (Saeidi et al., 2018; Sun et al., 2022; Kazemi et al., 2023), conditional reasoning abilities of LLMs remain unknown.
@@
Figure 1: Code prompting converts a natural language problem into a code prompt and prompts a large language model with such code to generate a natural language answer.
@@
Recently, researchers have improved performance on reasoning tasks by combining LLMs with symbolic interpreters such as the Python runtime  (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023) or SATisfiability solvers  (Ye et al., 2023). Here, LLMs pretrained on code or a combination of text and code (henceforth text+code LLM) are used to convert the input task into a symbolic language (e.g., Python or SAT problems), which is then fed into an external interpreter to make use of its symbolic execution. However, offloading the final reasoning task to an external execution module does not help with, or sometimes confounds, our understanding of the reasoning abilities of the model because part of the performance gain may come from the correct execution of the external interpreter. In particular, the fundamental questions of what contributes to the reasoning abilities and how reasoning abilities are triggered are still open questions needing further understanding. Nevertheless, pretraining on code is considered an important component that contributes to and explains the improved reasoning ability of LLMs. State-of-the-art LLMs such as GPT 3.5 (Kojima et al., 2022), GPT 4 (OpenAI, 2023), PaLM (Chowdhery et al., 2023), and Mistral 7B (Jiang et al., 2023) have been pretrained on both text and code and have demonstrated considerable boosts in many reasoning benchmarks.
@@
In this work, we investigate which aspects of code prompts, and in which way, elicit conditional reasoning abilities of GPT 3.5, a text+code LLM, across two datasets: (1) ConditionalQA (Sun et al., 2022), a scenario-based question answering (QA) dataset and (2) BoardgameQA (Kazemi et al., 2023), a boardgame-based QA dataset with conflicting rules. To understand the true benefit of code as an intermediate representation, we devise a chain of prompts, code prompting, that transform a natural language (NL) task into code and prompt the LLM with the generated code to elicit reasoning abilities. An illustration is provided in Figure 1. This setup has multiple benefits. Firstly, we fully utilize LLM’s reasoning ability without offloading any subtask to an external interpreter. In this way, we can make a fair comparison between text and code prompts without external variables such as interpreters. Secondly, enforcing compilability and executability of the generated code is not required, resulting in fewer interpreter-driven errors.
@@
  • We conduct a comprehensive study to compare code prompts with text prompts, showing i) large performance gains between $2.6$ and $7.7$ points while ii) being more efficient w.r.t. the number of demonstrations.
@@
  • We conduct an extensive analysis of why code prompts efficiently elicit conditional reasoning abilities in text+code LLMs, showing that prompting with code results in largely improved variable state tracking.
@@
Figure 2: Code Prompts converts a natural language problem instance into code before solving it with a large language model.
2 Related Work
@@
Most works generating code to solve natural language tasks use an external symbolic interpreter to run such code. Chen et al. (2023) and Gao et al. (2023) showed consistent gains on mathematical problems, symbolic reasoning, and algorithmic problems by using LLMs aided by external code interpreters. Lyu et al. (2023) show further gains in multi-hop QA, planning, and relational inference. Instead of leveraging a code interpreter, Ye et al. (2023) used an automated theorem prover with declarative code and showed consistent gains w.r.t. imperative code-interpreter-aided LLMs on arithmetic reasoning, logical reasoning, symbolic reasoning, and regex synthesis tasks. Lastly, Pan et al. (2023) did not use any interpreter and instead created programs composed of multiple subroutines and used smaller specialized models to run them. In this way, they outperform text prompts on text LLMs for fact-checking tasks.
@@
All these works employ an external solver system to run the code; therefore, the LLM is not conducting part of the reasoning. However, some works (Madaan et al., 2022; Liu et al., 2023b) suggest that LLMs of code may possess superior reasoning abilities than LLMs of text (i.e., pretrained only on natural language text). Madaan et al. (2022) observed improved commonsense reasoning, while Liu et al. (2023b) showed superior causal reasoning. This last work is based on the intuition that the large amount of if statements in the pretraining corpus enhances the causal reasoning abilities because they represent explicit causal relations. They conducted experiments on abductive and counterfactual reasoning tasks and showed that translating NL problems into code and then generating functions that return the answer to the problem with a code LLM outperforms prompting the same NL task in a text LLM. However, most popular LLMs are trained on text and code (e.g., GPT 3.5; Kojima et al. 2022, GPT 4; OpenAI 2023) and it remains unknown whether the reported performance gains come from the prompt format, the abilities of the LLM, or a combination of both. Therefore, it remains unclear whether code prompts also outperform text prompts in text+code models and, if so, why code prompts elicit reasoning abilities. In our work, we aim to answer all these questions.
@@
As far as we know, only the concurrent work of (Hussain et al., 2023) investigates the conditional reasoning abilities of LLMs. However, they only analyze the abilities of text-only LLMs after training them on ConditionalQA (Sun et al., 2022). Instead, we investigate how to use prompts to elicit conditional reasoning in text+code LLMs.
@@ 3.1 Definition
We hypothesize that querying text+code LLMs with instances translated to (pseudo)-code will result in improved conditional reasoning capabilities. Our hypothesis is motivated by previous results showing that program-aided LMs exhibit superior results on reasoning tasks than regular text prompts (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023; Ye et al., 2023). Thus, it comes naturally that prompting text+code models with instances translated to code could cause the underlying model to exhibit superior reasoning abilities when compared to text prompts. We formally define our hypothesis as follows:
@@
where $p$ is a problem instance from the set of problems $P$, $\sigma$ is an evaluation function that returns the quality of an answer for a given problem, and $T$ and $C$ are functions that create a natural language (Text) prompt and Code prompt, respectively, for the same given problem.
We define code prompts as prompts that model a natural language (NL) problem with code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. To solve an NL task with code prompts, we define a chain of prompts that i) transforms the NL text into code and ii) prompts using this code to generate the answer in natural language. Figure 1 illustrates this pipeline.
The prompt that transforms the NL problem to code is comprised of code that closely follows the original NL text. In particular, it creates variables for the key entities in the question and documents and if blocks for the documents, representing their conditional statements. Figure 2 illustrates this transformation.
@@ 3.2 Coding Features
To generate code as close as possible to the NL text, we use a programming language based on a simplification of Python. We only use boolean variables or variables that contain lists of strings. Variables follow the snake case naming convention. We also employ if statements to model conditional reasoning, but we do not use loops, functions, or classes. We create a code comment with the original NL text for each input sentence, and right after the code comment, we generate the code that represents the semantics of that sentence. However, we do not enforce that the generated code should be a runnable Python script.
@@ 4.1 Datasets
Throughout our experiments, we focus on two conditional question-answering QA datasets: ConditionalQA (CondQA; Sun et al., 2022) and BoardgameQA (BGQA; Kazemi et al., 2023). Solving these datasets requires advanced conditional and compositional reasoning capabilities.
ConditionalQA is a QA dataset where the answers are applicable under specific scenarios (i.e., conditional answers). Therefore, along with each question, the dataset provides a scenario that describes the background of the person posing such a question. Questions require multi-hop, compositional, and conditional logic over documents about public policies (e.g., the eligibility for a subsidy). Answers can be a span of the document, yes, and no. We use an oracle retriever to select the relevant passages to the question so that we can isolate the analysis of conditional reasoning abilities in LLMs from the retrieval component. The expected output is a chain of thought (CoT; Wei et al. 2022) followed by the final answer. To create the CoT, we use the annotated evidence sentences.
BoardgameQA is a dataset that evaluates the ability to reason with contradictory information guided by preferences. For example, given a question about traveling abroad, information found online about regulations can be contradictory because rules may change over time. Answering questions in this dataset requires complex multi-hop reasoning with conditional, deductive, and compositional abilities. The domain of the problems is board games, which allows us to analyze the conditional reasoning abilities in a completely different domain from CondQA. BGQA is divided into multiple partitions focusing on different characteristics, such as the depth of the reasoning tree, the need for external information, etc. We focus on the main partition and its subpartitions (i.e., BGQA-1, BGQA-2, BGQA-3), where the number refers to the number of reasoning hops required to answer the question. This dataset also includes annotated chain-of-thoughts (CoT); therefore, we use their annotated input (“example”) as the input prompt and their annotated CoT (“proof”) as the expected output.
@@
We include more details about the datasets in Appendix A, a formal definition of the prompts in Appendix B, and examples in Appendix G.
4.2 LLM Setup.
We perform our study using gpt-35-turbo-0613-16k through the Azure OpenAI service. All of our prompting methods are implemented using the Langchain library. [^2] https://github.com/langchain-ai/langchain We set the decoding temperature to zero and use greedy sampling to make the outputs deterministic. We execute our prompts with in-context learning. However, the amount of demonstrations is constrained by their long length and the context window of the model, especially in code prompts, which are longer than text prompts because they include the original natural language text and the corresponding code. Specifically, we used between three and nine demonstrations except for code prompts in BGQA-2 and BGQA-3 where the maximum possible is 6. In other words, we provided from one to three demonstrations per class. For each experiment, we use a random sample from the training set as demonstrations. A detailed analysis of the impact of the number of demonstrations in the performance of the prompts is provided in Section 5.4. We also report the costs of our prompts in Appendix C.
4.3 Evaluation.
@@
We follow the evaluation metrics used in the original datasets. For CondQA, we report the F1 token overlap between the predicted answer and the label, while for BGQA, we report accuracy. We run the main experiments multiple times with different random seeds: two times on BGQA and three times on CondQA. We report the average and standard deviation performance across these runs. For the subsequent analyses of code prompts, we run each experiment once due to the inference costs.
Prompt Method CondQA BGQA-1 BGQA-2 BGQA-3
Text Prompt $56.32\pm 1.06$ $62.80\pm 0.28$ $55.00\pm 0.56$ $47.59\pm 0.00$
Code Prompt 59.17 $\pm$ 0.74 69.40 $\pm$ 1.13 62.00 $\pm$ 1.70 55.30 $\pm$ 0.42
Atomic Statements 55.94 57.80 47.80 35.60
Code $\rightarrow$ NL 55.42 63.00 55.60 48.00
Table 1: Comparison of text prompt and code prompts on the validation set of each dataset. Last two rows analyse the impact of the input format on text prompts.

### 25 Feb 2024 v2

@@ Abstract
Reasoning is a fundamental component of language understanding. Recent prompting techniques, such as chain of thought, have consistently improved LLMs’ performance on various reasoning tasks. Nevertheless, there is still little understanding of what triggers reasoning abilities in LLMs in the inference stage. In this paper, we introduce code prompting, a chain of prompts that transforms a natural language problem into code and directly prompts the LLM using the generated code without resorting to external code execution. We hypothesize that code prompts can elicit certain reasoning capabilities of LLMs trained on text and code and utilize the proposed method to improve conditional reasoning, the ability to infer different conclusions depending on the fulfillment of certain conditions. We find that code prompting exhibits a high-performance boost for multiple LLMs (up to $22.52$ percentage points on GPT 3.5, $7.75$ on Mixtral, and $16.78$ on Mistral) across multiple conditional reasoning datasets. We then conduct comprehensive experiments to understand how code prompts trigger reasoning abilities and which capabilities are elicited in the underlying models. Our analysis of GPT 3.5 reveals that the code formatting of the input problem is essential for performance improvement. Furthermore, code prompts improve sample efficiency of in-context learning and facilitate state tracking of variables or entities. [^1] Our code and prompts are available at this URL.
@@ 1 Introduction
Reasoning is a fundamental component of both human and artificial intelligence (AI) and the backbone of many NLP tasks, such as question answering and textual entailment. Recently, intensive studies have been performed on mathematical (Patel et al., 2021; Chen et al., 2021; Cobbe et al., 2021), logical (Liu et al., 2020, 2023a; Sinha et al., 2019), and commonsense reasoning (Madaan et al., 2022; Liu et al., 2022a, b; Wang et al., 2023). Conditional reasoning, a primary yet complex reasoning ability that draws alternative conclusions depending on the fulfillment of certain conditions, remains understudied. These conditions are stated in the text, making the problem self-contained, which allows us to study the semantic inferencing capabilities of the underlying model, i.e., identifying relevant premises and ascertaining the presence of total evidence Nolt et al. (1988); Cabria and Magnini (2014) without the requirement for, and confounding effects of external knowledge. Conditional reasoning is also a fundamental form of logical reasoning useful in many practical scenarios, such as answering real-world questions about the eligibility for a visa or a loan. Despite the recent introduction of some benchmarks (Saeidi et al., 2018; Sun et al., 2022; Kazemi et al., 2023), conditional reasoning abilities of LLMs remain unknown.
@@
Figure 1: Code prompting converts a natural language problem into a code prompt and prompts a large language model with such code to generate an answer.
@@
Recently, researchers have improved performance on reasoning tasks by combining LLMs with symbolic interpreters such as the Python runtime  (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023) or SATisfiability solvers  (Ye et al., 2023). Here, LLMs pretrained on code or a combination of text and code (henceforth text+code LLM) are used to convert the input task into a symbolic language (e.g., Python or SAT problems), which is then fed into an external interpreter to make use of its symbolic execution. In such a setup, the LLM is mainly used for the linguistic dimension of reasoning (i.e., identifying the premises from text) and planning how to solve the problem, while the logical reasoning component (i.e., solving the actual logical problem) is offloaded to an external execution module, confounding our understanding of the reasoning abilities of the underlying model. In particular, the fundamental questions of what contributes to the reasoning abilities and how reasoning abilities are triggered remain open. Nevertheless, pretraining on code is considered an important component that contributes to and explains the improved reasoning ability of LLMs. State-of-the-art LLMs such as GPT 3.5 (Kojima et al., 2022), GPT 4 (OpenAI, 2023), Mixtral (Jiang et al., 2024), and Mistral 7B (Jiang et al., 2023) have been pretrained on both text and code and have demonstrated considerable boosts in many reasoning benchmarks.
@@
In this work, we focus on improving the logical reasoning dimension of semantic inference and investigate which aspects of code prompts, and in which way, elicit conditional reasoning abilities of several text+code LLMs (GPT 3.5, Mixtral, and Mistral), across three datasets: (1) ConditionalQA (Sun et al., 2022), (2) BoardgameQA (Kazemi et al., 2023) and (3) ShARC (Saeidi et al., 2018). To understand the benefit of code as an intermediate representation, we devise a chain of prompts, code prompting, that transform a natural language (NL) task into code and directly prompts the LLM with the generated code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. An illustration is provided in Figure 1. This setup has multiple benefits. Firstly, we fully utilize the LLM’s reasoning abilities without offloading any subtask to an external interpreter. In this way, we can conduct a fair comparison between text and code prompts without external variables such as interpreters. Secondly, enforcing compilability and executability of the generated code is not required, resulting in fewer interpreter-driven errors.
@@
  • We conduct a comprehensive study to compare code prompts with text prompts, showing (i) large performance gains on the three LLMs (up to $22.52$ points for GPT3.5, up to $7.75$ for Mixtral, and up to $16.78$ for Mistral), while (ii) being more efficient with regard to the number of demonstrations.
@@
  • We conduct an extensive analysis to understand why code prompts efficiently elicit conditional reasoning abilities, and show that prompting with code results in largely improved variable state tracking.
@@
Figure 2: Code prompting converts natural language descriptions into code to be solved with a large language model. The figure shows a transformed instance from the ConditionalQA dataset.
2 Background and Related Work
Semantic Inference.
It is important to understand the framework within which we conduct our experiments. Question answering can be seen as a chain of semantic inferences. Semantic inference can be decomposed into (1) the linguistic dimension, which identifies implications from natural language text and (2) the logical dimension, where reasoning, deductive, inductive or abductive, is conducted given the identified premises  (Cabria and Magnini, 2014). The logical dimension resides on four major criteria which a reasoning model should validate Nolt et al. (1988): (a) the truth of premises, (b) validity and inductive probability of premises, (c) relevance of premises to the conclusion, and (d) requirement of total evidence. In the scope of our work, we focus on improving deductive reasoning along the logical dimension. We assume the truthfulness of information presented within the datasets, making the task of the analyzed LLM to (c) identify relevant premises and (d) ascertain whether total required evidence is present, which is the focus of conditional reasoning.
Code LLMs.
@@ 2 Background and Related Work
Most works that generate code to solve natural language tasks use an external symbolic interpreter to run the resulting code. Chen et al. (2023) and Gao et al. (2023) showed consistent gains on mathematical problems, symbolic reasoning, and algorithmic problems by using LLMs aided by external code interpreters. Lyu et al. (2023) further showed the improvement in multi-hop QA, planning, and relational inference. In contrast, Ye et al. (2023) used an external automated theorem prover with declarative code and showed consistent gains w.r.t. imperative code-interpreter-aided LLMs on arithmetic reasoning, logical reasoning, symbolic reasoning, and regex synthesis tasks. Lastly, Pan et al. (2023) did not use any interpreter and instead created programs composed of multiple subroutines and used smaller specialized models to run them. In this way, they outperform text prompts on text LLMs for fact-checking tasks.
@@
All these works employ an external solver system to run the code; therefore, the LLM is not conducting part of the reasoning. Some works (Madaan et al., 2022; Liu et al., 2023b) suggest that LLMs trained on code may possess superior reasoning abilities than LLMs trained only on natural language text. Madaan et al. (2022) observed improved commonsense reasoning, while Liu et al. (2023b) report superior causal reasoning. The latter work is based on the intuition that large amounts of if statements in the pretraining corpus enhance causal reasoning abilities because they represent explicit causal relations. They conducted experiments on abductive and counterfactual reasoning tasks and showed that translating NL problems into code and then generating functions that return the answer to the problem with a code LLM outperforms prompting the same NL task in a text LLM. However, most popular LLMs are trained on text and code (e.g., GPT 3.5; Kojima et al. 2022, GPT 4; OpenAI 2023). Therefore, it remains unclear whether code prompts also outperform text prompts in text+code models and, if so, why code prompts elicit reasoning abilities. In our work, we aim to answer these questions.
@@
To the best of our knowledge, only the concurrent work of Hussain et al. (2023) investigates the conditional reasoning abilities of LLMs. However, they only analyze the abilities of text-only LLMs after training them on ConditionalQA (Sun et al., 2022). Instead, we investigate how to use prompts to elicit conditional reasoning in text+code LLMs.
@@ 3.1 Definition
We hypothesize that querying text+code LLMs with instances translated to code will result in improved conditional reasoning capabilities. Our hypothesis is motivated by prior works showing that program-aided LMs exhibit superior results on reasoning tasks than regular text prompts (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023; Ye et al., 2023). Thus, it comes naturally that prompting text+code LLMs with instances translated to code could cause the underlying model to exhibit superior reasoning abilities when compared to text prompts. We formalize our hypothesis as follows:
@@
where $p$ is a problem instance from the set of problems $P$, $\sigma$ is an evaluation function that returns the quality of an answer for a given problem, and $T$ and $C$ are functions that create a natural language Text prompt and Code prompt, respectively, for the same given problem.
We define code prompts as prompts that model a natural language (NL) problem with code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. To solve an NL task with code prompts, we define a chain of prompts that i) transform the NL text into code, and ii) use this code to generate the answer in natural language. Figure 1 illustrates this pipeline.
The prompt transforming the NL problem into code is composed of code that closely follows the original NL text. In particular, it creates variables for key entities in the question and documents and if blocks for each conditional statement in the documents. Figure 2 exemplifies this transformation.
@@ 3.2 Coding Features
To generate code as close as possible to the NL text, we use a programming language based on a simplification of Python. We only use boolean variables or variables that contain lists of strings. Variables follow the snake case naming convention. We also employ if statements to model conditional reasoning, but we do not use loops, functions, or classes. We create a code comment with the original NL text for each input sentence, and right after the code comment, we generate the code that represents the semantics of that sentence. However, we do not enforce the generated code to be a runnable script.
@@ 4.1 Datasets
Throughout our experiments, we use three question-answering (QA) datasets for conditional reasoning: ConditionalQA (CondQA; Sun et al., 2022), a scenario-based question answering (QA) dataset, BoardgameQA (BGQA; Kazemi et al., 2023), a boardgame-base QA dataset with conflicting rules, and ShARC (Saeidi et al., 2018), a conversational QA dataset with natural language rules. Solving these datasets requires advanced conditional and compositional reasoning capabilities.
We focus on the QA task of CondQA. For BGQA, we focus on the main partition, which includes three subsets BGQA-1, BGQA-2, and BGQA-3, where the number indicates the reasoning hops needed to answer. Lastly, while ShARC encompasses dialogue generation, we aim to evaluate specific capabilities unrelated to conversational flow. Therefore, we isolated the QA pairs from the provided dialogues, resulting in a dataset where the model has to answer yes, no, or not enough information. [^2] In the full task, not enough information would trigger another step in a pipeline to generate a follow-up question.
@@
We include more details about the datasets in Appendix A, a formal definition of the prompts in Appendix B, and examples in Appendix K.
4.2 LLM Setup
We perform our study using gpt-35-turbo [^3] https://platform.openai.com/docs/models/gpt-3-5-turbo through the Azure OpenAI service, Mixtral 7x8B (Jiang et al., 2024), and Mistral 7B (Jiang et al., 2023) with a Nvidia A100. The use of these models allows us to investigate whether our approach holds across different model sizes. In particular, Mistral 7B contains 7 billion parameters, and Mixtral 7x8B 46.7 billion. All of our prompting methods are implemented using the Langchain library. [^4] https://github.com/langchain-ai/langchain We set the decoding temperature to zero and use greedy sampling to make the outputs deterministic. We execute our prompts with in-context learning and provide one demonstration per class. For each experiment, we use a random sample from the training set as demonstrations. The LLM generating the code for code prompts is the same one as the one running the code to generate the final answer. We evaluate each model and prompt in the dev set of each dataset with two random seeds. Since the demonstrations are selected randomly, the seed determines them. The seed that yields the best performance on the dev set is then used for the final evaluation on the test set. We provide more details on the models and the costs in Appendix C and D.
4.3 Evaluation
@@
We follow the evaluation metrics used in the original datasets. For CondQA, we report the F1 token overlap between the predicted answer and the label, while for BGQA and ShARC, we report the macro F1 score. We run the main experiments two times with different random seeds (00 and $1$). We report the average and standard deviation performance across these runs. For the subsequent analyses of code prompts, we run each experiment once only on GPT 3.5 due to the inference costs.
Model Prompt CondQA ShARC BGQA-1 BGQA-2 BGQA-3 $\Delta$CP
Test Set
GPT 3.5 Text $58.70$ $\mathbf{63.78}$ $51.15$ $37.42$ $27.77$ 9.17
Code $\mathbf{60.60}$ $59.54$ $\mathbf{58.67}$ $\mathbf{55.56}$ $\mathbf{50.29}$
Mixtral Text $\mathbf{48.17}$ $60.29$ $\mathbf{56.38}$ $39.64$ $30.15$ 3.66
Code $44.73$ $\mathbf{62.77}$ $53.33$ $\mathbf{47.39}$ $\mathbf{44.72}$
Mistral Text $\mathbf{35.74}$ $37.70$ $47.40$ $48.78$ $47.86$ 4.83
Code $33.28$ $\mathbf{54.48}$ $\mathbf{53.80}$ $\mathbf{51.27}$ $\mathbf{48.79}$
Dev Set
GPT 3.5 Text $56.54\pm 0.08$ $\mathbf{63.91}\pm\mathbf{0.95}$ $53.16\pm 1.67$ $33.71\pm 10.37$ $31.5\pm 13.39$ 9.79
Code $\mathbf{57.64\pm 1.42}$ $58.13\pm 1.62$ $\mathbf{68.60\pm 1.09}$ $\mathbf{55.85\pm 4.06}$ $\mathbf{47.57\pm 2.68}$
Mixtral Text $\mathbf{46.60}\pm\mathbf{0.99}$ $53.55\pm 1.58$ $58.31\pm 1.77$ $36.77\pm 0.09$ $32.06\pm 1.79$ 2.76
Code $40.88\pm 1.84$ $\mathbf{58.03}\pm\mathbf{2.81}$ $57.94\pm 5.52$ $\mathbf{45.32}\pm\mathbf{0.54}$ $\mathbf{38.90}\pm\mathbf{7.33}$
Mistral Text $28.84\pm 0.02$ $37.30\pm 1.69$ $47.61\pm 0.92$ $47.29\pm 1.97$ $46.56\pm 2.92$ 5.05
Code $28.26\pm 10.03$ $\mathbf{52.90}\pm\mathbf{1.08}$ $\mathbf{52.21}\pm\mathbf{0.95}$ $\mathbf{54.27}\pm\mathbf{1.42}$ $45.22\pm 10.75$
Table 1: Comparison (F1 score) of text prompt and code prompts. All results use one demonstration per class.    $\Delta$CP = Code Prompt - Text Prompt, i.e., the average performance gain from code prompts across all datasets.